# The ultimate issue error: mistaking parameters for hypotheses

### v1

@@ Discussion
Therefore, to maximize the WoE an experiment can provide, researchers should strive to keep the nominal false positive rate below $\alpha=0.05$. This should come as no surprise, since minimising bias, removing confounding, and ruling out alternative explanations are effective ways to improve an experiment.
However, when providing evidence for the null hypothesis, it is better invest effort to minimise false negatives (numerator of Equation 5), which can be verified with similar calculations.
$P(\text{Model})$ and $P(\text{Hypothesis})$
As an alternative to hypothesis testing, model comparison is another method of inference (Maxwell and Delaney 2004; Judd, McClelland, and Ryan 2009), which can be classified into two types. The first case is when two models are nested. This situation arises when one model restricts one or more parameters to specific values (usually the null value of zero or one), and the other model allows these parameters to vary and to be estimated from the data. A comparison is then made between the restricted and unrestricted models to determine if the additional flexibility of the unrestricted model produces a better fit to the data, after penalising it for its additional flexibility. This is simply another method of parameter testing since it assesses whether a parameter equals some null value or not, and the p-values will often be the same. In this case, all of the above comments are applicable.
The other type of comparison occurs when the models are not nested, that is, when one model cannot be reduced to the other by fixing parameter. An example is these two nonlinear models: $\theta_{2}(1-e^{-\theta_{1}x})$ and $\theta_{1}x/(\theta_{2}+x)$. They have qualitatively similar behaviour in that $y=0$ when $x=0$, and the functions rise to an upper asymptote as $x$ gets larger. The models provide two different stories about how the data were generated, much like the prosecution and defence provide different stories about how the evidence arose. In such cases model testing and hypothesis testing are more alike, especially when the models represent mechanistic or causal relationships and not just different empirical models. Nevertheless, background assumptions and context are still necessary (e.g. were the measurement methods sound, was the design appropriate, do the authors own shares of a company that benefits only if one model is better, etc.).
@@ Conclusion
We have argued that in most biological and social science research the probability of a parameter cannot equal the probability of a hypothesis. First, qualitative background and contextual information is always needed to interpret the parameter value in light of the hypothesis. Second, multiple parameters associated with multiple outcomes that test the same hypothesis are rarely equal and may even conflict. Third, the value of a parameter (and its corresponding p-value or Bayes factor) often depends on unknown and unmeasured characteristics of the population or experiment.
@@
The proposed approach relies on subjective judgement to weigh evidence and interpret results, and interpretations may differ among researchers. However, this is only a formalization of the process most scientists use when reading a manuscript. A formalisation can help identify where the problem(s) are, and if one is willing to put numbers to actual power and the probability of a false positive, a quantitative estimate of the support for one hypothesis over another can be derived. In addition, formalisation can help resolve disagreements by breaking a complex inferential task into more manageable components that are evaluated separately, and then combined for a final WoE. Using this approach, for example, one researcher might conclude that $H_{1}$ has little support due to the authors’ conflict of interest, whereas another might conclude that the lack of blinding and randomisation undermines the study. This process is no different from peer review or an editorial decision to accept or reject a manuscript based on the quality of the research.
WoE provides a structured approach, but there is a risk that its subjectivity and flexibility could be exploited to justify biased interpretations, particularly in contentious or high-stakes research areas. For instance, senior scientists entrenched in the dominant paradigm could use the WoE approach to dismiss novel and promising research avenues. The WoE method, however, promotes a more comprehensive and transparent integration of evidence, which enables critical scrutiny of both data and underlying assumptions, and which can ultimately lead to more robust and credible conclusions.

### v2

@@ Discussion
Therefore, to maximize the WoE an experiment can provide, researchers should strive to keep the nominal false positive rate below $\alpha=0.05$. This should come as no surprise, since minimising bias, removing confounding, and ruling out alternative explanations are effective ways to improve an experiment. However, when providing evidence for the null hypothesis, it is better to minimise false negatives (numerator of Equation 5), which can be verified with similar calculations.
Evaluating manuscripts and published evidence
The proposed approach relies on subjective judgement to weigh evidence and interpret results, and interpretations may differ among researchers. However, this is only a formalization of the process most scientists use when reading a manuscript. Formalisation can help identify where the problem(s) are, and if one is willing to put numbers to actual power and the probability of a false positive, a quantitative estimate of the support for one hypothesis over another can be derived. This process is no different from what peer reviewers and editors informally do when deciding to accept or reject a manuscript based on the quality of the research. It is an empirical question whether such an approach would improve peer-review and the editorial assessment of manuscripts, perhaps with the aid of a checklist to standardise the criteria considered.
Similarly, published papers can be evaluated with a WoE approach, as was done for the vitamin D and COVID-19 example. Papers such as Hastie et al. (2020) provide nearly zero information for a vitamin D/COVID-19 association and can therefore be ignored, and certainly not included in meta-analyses, where they can bias the results.
@@ Conclusion
We have argued that in most biological and social science research the probability of a parameter cannot equal the probability of a hypothesis. First, qualitative background and contextual information is always needed to interpret the parameter value in light of the hypothesis. Second, multiple parameters associated with multiple outcomes that test the same hypothesis are rarely equal and may even conflict. Third, the value of a parameter (and its corresponding p-value or Bayes factor) often depends on unknown and unmeasured characteristics of the population or experiment as well as the statistical model.
@@
WoE provides a structured approach, but there is a risk that its subjectivity and flexibility could be exploited to justify biased interpretations, particularly in contentious or high-stakes research areas. For instance, senior scientists entrenched in the dominant paradigm could use the WoE approach to dismiss novel and promising research avenues. The WoE method, however, promotes a more comprehensive and transparent evaluation of evidence, which enables critical scrutiny of both data and underlying assumptions, ultimately leading to more robust and credible conclusions.